OpenStudy (anonymous):

y/-2+5=13 ~solve for y ~ I NEED THIS ASAP <3

OpenStudy (mertsj):

What is -2+5?

OpenStudy (anonymous):

3

OpenStudy (mertsj):

What number, divided by 3, is equal to 13?

OpenStudy (anonymous):

idk.., i dont understand

OpenStudy (mertsj):

\[\frac{y}{-2+5}=13\]

OpenStudy (mertsj):

Is that your problem?

OpenStudy (mertsj):

Or is it this: \[\frac{y}{-2}+5=13\]

OpenStudy (anonymous):

yea except only -2 is under y not the +5

OpenStudy (anonymous):

yeaa! that one. : )

OpenStudy (mertsj):

1. Subtract 5 from both sides 2. Multiply both sides by -2

OpenStudy (anonymous):

-16.?

OpenStudy (anonymous):

OMG YESS Graciass.!! : D

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

yw